Kgabo completed 2/5 of a job. The next day he completed 3/8 of the remaining job. The part of the job that is still left, given as a fraction is?


1
Expert's answer
2022-03-03T04:48:36-0500

Let the total job be 1. So Kgabo completed 2/5 of a job and "1-\dfrac25 = \dfrac35" of a job was left. The next day he completed 3/8 from the left job or "\dfrac38\cdot\dfrac35 = \dfrac{9}{40}" of a total job, so it is left only "1-\dfrac25 - \dfrac{9}{40} = \dfrac35 - \dfrac{9}{40} = \dfrac{24}{40} - \dfrac{9}{40} = \dfrac{15}{40} = \dfrac{3}{8}" of a total job.